# CLEAR: Towards Contextual LLM-Empowered Privacy Policy Analysis and Risk Generation for Large Language Model Applications

### v2

@@ 8.3.1. Study design
Although the results of the evaluation study suggest the high usability and usefulness of our system, they are limited to a short-term study in a lab setting. A future deployment study is needed to validate these findings in long-term and real-world settings. Our study is an initial effort to enable users to investigate relevant privacy policies and potential privacy risks when using LLMs. In the future, we plan to collect additional behavioral data when users interact with CLEAR in the deployment. This will help identify the key factors and issues that influence users’ privacy awareness when using our system. We will also conduct pre-tests and post-tests with users to assess their acquisition of privacy knowledge and perform longitudinal deployment studies to measure changes in user behavior over time.
@@ 8.3.2. System design
Our current CLEAR prototype serves primarily as a proof of concept to facilitate the experiment discussed in this paper. At present, it is compatible only with ChatGPT and the Gemini plugin for Gmail in order to demonstrate the feasibility of CLEAR in two primary use cases (standalone LLM conversational interfaces and the LLM add-ons for GUI applications). While ChatGPT and Gmail are the most popular applications in their category, with ChatGPT having a 60% market share (Westfall, 2023) and Gmail holding a 53% share of the US email market [^12] https://techreport.com/statistics/software-web/gmail-statistics/, there are other popular LLMs (e.g., Claude) and applications (e.g., Microsoft Copilot) that the current CLEAR tool does not support. Our system also has limited capabilities in identifying additional modalities of sensitive information (e.g., through voice interaction) and handling sensitive information contained in files such as images and PDFs uploaded to LLMs. Moreover, while we effectively use the Microsoft Presidio library to detect emails, contacts, and locations, CLEAR currently lacks the capability to promptly identify other types of personally identifiable information, such as financial and health data. To enhance CLEAR’s functionality, we intend to expand support to additional LLM platforms and improve its ability to detect a broader range of sensitive information types and modalities.

### v3

@@ 8.3.1. Study design
Although the results of the evaluation study suggest the high usability and usefulness of our system, they are limited to a short-term study in a lab setting. A future longitudinal deployment study is needed to validate these findings in long-term and real-world settings. Our study is an initial effort to enable users to investigate relevant privacy policies and potential privacy risks when using LLMs. In the future, we will recruit a larger and more representative participant group through online recruitment platforms (e.g., Prolific), ensuring diversity across demographics, levels of technical proficiency, and privacy awareness backgrounds. Participants will install the CLEAR extension and interact with LLM conversational interfaces or LLM-based add-ons in the real context of use for 2 weeks. We plan to conduct pre-tests and post-tests with users to assess their acquisition of privacy knowledge and collect additional behavioral data when users interact with CLEAR in the deployment (similar to (Story et al., 2021, 2022)). This future longitudinal study will allow us to evaluate the ecological validity of CLEAR and the impact of its use on the user’s real privacy behaviors.
@@ 8.3.2. System design
Our current CLEAR prototype serves primarily as a proof of concept to facilitate the experiment discussed in this paper. At present, it is compatible only with ChatGPT and the Gemini plugin for Gmail in order to demonstrate the feasibility of CLEAR in two primary use cases (standalone LLM conversational interfaces and the LLM add-ons for GUI applications). While ChatGPT and Gmail are the most popular applications in their category, with ChatGPT having a 60% market share (Westfall, 2023) and Gmail holding a 53% share of the US email market [^12] https://techreport.com/statistics/software-web/gmail-statistics/, there are other popular LLMs (e.g., Claude) and applications (e.g., Microsoft Copilot) that the current CLEAR tool does not support. However, the key features of CLEAR do not rely on any specific technicalities of ChatGPT or Gemini for Gmail, and it is straightforward to extend CLEAR to support other chat-based LLM interfaces (e.g., Claude, web UIs for LLaMA and Qwen) and web-based LLM add-ons. To expand its scalability, we open-sourced CLEAR  [^13] https://github.com/CRChenND/CLEAR. CLEAR’s extensible architecture will make it easier to integrate CLEAR with other popular LLMs’ GUI in the future.
Our system also has limited capabilities in identifying additional modalities of sensitive information (e.g., through voice interaction) and handling sensitive information contained in files such as images and PDFs uploaded to LLMs. Moreover, while we effectively use the Microsoft Presidio library to detect emails, contacts, and locations, CLEAR currently lacks the capability to promptly identify other types of personally identifiable information, such as financial and health data. To enhance CLEAR’s functionality, we plan to investigate using small-scale, open-source language models to identify sensitive information on the edge. Recent advancements (Zhou et al., 2024; Li et al., 2024b) allow these models to run locally on the client side, allowing for more accurate detection of sensitive information across a broader range of types, while keeping user data on-device and mitigating privacy risks.
Currently, CLEAR uses text to summarize privacy policies and generate privacy risks as an initial effort to enable users to investigate relevant privacy policies and potential risks when using LLMs. For users with limited privacy literacy or technical expertise, textual descriptions may not be the most efficient way to communicate the implications of privacy practices or the significance of identified risks. With the development of multimodal LLMs, there is an opportunity to overcome these limitations by incorporating more interactive and accessible explanations. Future work could focus on integrating multimodal features into CLEAR, such as interactive tutorials or visual summaries that highlight key aspects of privacy policies and risks. For example, users could explore how their data might be processed or shared through animations or flowcharts that depict data flows. Integrating multimodal LLMs into CLEAR has the potential to transform how privacy policies and risks are communicated, making them more transparent, engaging, and comprehensible for a wider audience.
@@ 9. Conclusion
Acknowledgements.
This work was supported in part by a Google Research Scholar Award, a Notre Dame-IBM Technology Ethics Lab Award, a Google PSS Privacy Research Award, NSF Grant CNS-2426395, NSF Grant CNS-2426397, and NSF Grant CNS-2232653. C. Chen and Y. Ye’s work was partially supported by the NSF under grants IIS-2321504, IIS-2334193, IIS-2217239, CNS-2426514, CNS-2203261, and CMMI-2146076. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of the sponsors. We would like to thank Allen Yilun Lin for useful discussions.